# Predictive Non-linear Dynamics via Neural Networks and Recurrence Plots

### v1

@@ II.3 CNN and metrics
Each network was trained independently in two different RP train sets: the first set containing data from the Logistic map generated with parameters $r\in[3.0,4.0]$, and the second with $r$ immediately after the onset of chaos; while for the Standard map, $k\in[0.1,0.971635)$ and a larger range $k\in[0.1,2.0]$ were used to generate the data sets for this dynamic. The amount of samples, originated from the same sets of ICs, remained constant for the analysis focused on the onset of chaos and $k<k_{c}\approx 0.971635$, resulting in larger training sets for these regions where the chaotic behavior is predominant.
@@
The evaluation of the performance of the trained models used the following metrics: the weighted $f_{1}$ score was employed to assess the accuracy of predictions across the defined intervals of $r$ to which the testing RP samples belonged in the Logistic map. For the regression tasks, a point-wise error, based on the absolute differences between the actual and predicted parameter values, was used to evaluate the estimated $r$ and $k$ values for the Logistic and Standard maps, respectively.
Additionally, considering an acceptable margin $m$, a given prediction can be categorized as follows:
@@
where $y_{i}$ and $\hat{y_{i}}$ represent the actual and predicted parameter values for sample $i$, respectively. The Heaviside step function ensures that if $|y_{i}-\hat{y_{i}}|<m$, $C_{i}=1$, indicating that the predicted value is within the acceptable margin $m$ from the actual value; otherwise, $C_{i}=0$, representing an incorrect prediction.
The total number of predictions within the $m$-tolerance is then given by $C(m)=\sum_{i=1}^{N}C_{i}(m)$, where $N$ is the total number of test samples. Furthermore, the percentage of correct predictions is defined as
@@
where, in practical terms, if $m=0.01$, the percentage $c(0.01)$ expresses how accurately the NN can predict the parameter value up to the third decimal place.
@@ III Results and discussions
The first set of results comes from applying the aforementioned methodology to the Logistic map. $4\times 10^{4}$ RPs, representing distinct non-linear dynamical behaviors from Eq. (6), were constructed considering fixed $RR=20\%$ and $T=100$. The task of the trained NNs is to predict both the value and range of the characteristic control parameter $r$, respectively through a regression and classification task. Table 1 presents the outcome of the models for the hold-out test samples.
Table 1: Accuracy presented for the classification task on the logistic map, employing $f_{1}$ score as the metric; and the regression task, using $RMSE$, both for the test set, for each interval of $r$ as described in Sec. II.2.1.
@@
(A3) 0.0069 100% 82.35% 0.971 17
(B) 0.0459 95.07% 24.34% 0.965 608
(C) 0.0722 86.93% 16.00% 0.947 375
weight. avg. 0.0551 - - 0.989 1000
Figure 6: Predicted values ($\hat{r}$) compared to actual parameter ($r$) values of the test set from the Logistic Map in the regression task. The colour axis illustrates the dispersion on the absolute difference $|r-\hat{r}|$ in both cases: (a) $r\in[3.0,4.0]$ and; (b) $r\in[3.56,4.0]$.
Notice that the weighted accuracy obtained, shown in column “$f_{1}$ score” for training the NN almost exclusively on the chaotic region is higher than currently reported in literature [9, 10], and also improved while comparing with RPs originated from $r\in[3.0,4.0]$. As for the regression results displayed on columns “RMSE, c(0.1) and c(0.01)”, the same improvement on characterizing chaotic dynamics is observed when focusing solely on the chaotic region, where at least $86.93\%$ of predicted parameters $\hat{r}$ are precise within two decimal places to the actual parameter $r$.
Figure 6 illustrates the regression results for the Logistic map. From Fig. 6 (a), is clear that the proposed NN is successfully predicting the parameter values, especially before the onset of chaos at $r\approx 3.57$. After that, the predictions weaken as the chaotic regime dominates the dynamics. To further investigate, the same amount of RP samples was used to train the NN, solely considering $r\in[3.56,4.0]$. The regression for this specific range is shown in Fig. 6 (b).
In comparison to (a), Fig. 6 (b) shows an inferior agreement between the predicted and actual parameter values. This is due to the fact that the chaotic regime may present many distinct recurrence patterns, as earlier evidenced by the RPs within this range in Fig. 2. These many faces of chaotic behavior are expected for a finite-time analysis, making the predictions more challenging, and consequently lowering the NN accuracy. Additionally, regions between $3.93.8<r<$ 3.9, where the agreement is even less accurate, correlate to the rise of periodic and quasi-periodic behaviors due to islands of stability of said region. Yet, as observed in Tab. 1, this region is successfully classified, presenting relatively high $f_{1}$ score.
Figure 7: Predicted values ($\hat{k}$) compared to actual parameter ($k$) values of the test set from the Logistic map in the regression task. The color axis illustrates the dispersion on the absolute difference $|k-\hat{k}|$ in both cases: (a) $k\in[0.1,k_{c})$ and; (b) $r\in[0.1,2.0)$.
Table 2: RMSE and overall precision for the regression task for RPs generated from the Standard map.
@@
$[0.1,k_{c})$ 0.0273 99.4% 47.8%
$[0.1,2)$ 0.0742 85.0% 14.6%
Considering the Standard map, Tab. 2 summarizes the performance of the trained NN with the RPs from the different non-linear dynamical behaviors of Eqs. (7) and (8). Again, $4\times 10^{4}$ RPs were constructed considering $T=100$, but now $RR=10\%$. In contrast to the Logistic map, all RPs of the Standard map represent chaotic behaviors, specifically different finite-time trajectories in Hamiltonian chaos. In that sense, employing the process outlined in Sec. II.1 for a relatively high $RR$ may require longer computational simulations, so $RR=10\%$ was chosen. Additionally, since only chaotic behavior is analyzed, classification tasks were neglected.
It is important to emphasize that by selecting trajectories with specific ICs, namely $(q_{0},p_{0})=(\delta\times 10^{-8},\pi)$, where $\delta=2,4,6~{}\text{and}~{}8$, is possible to ensure chaotic behavior for $k>0.1$. As detailed in Sec. II.2, for $k\gtrsim 0.1$, a chaotic layer surrounds the vicinity of the unstable fixed point at $(0,\pi)$ and, as $k$ increases, the phase space is filled by the expanding chaotic sea. These chaotic trajectories are suitably represented by their respective RPs as depicted earlier by Fig. 4.
Regression results for the Standard map are shown in Fig. 7. Initially in (a), considering the interval $k\in[0.1,k_{c})$, the agreement between the predicted and actual parameter values is surprisingly notable for chaotic dynamics. A slight decrease in performance is observed in (b) for parameters beyond the critical value $k_{c}$. Although regions above $k>2.0$ were neglected, these results illustrate that particular finite-time chaotic behaviors, associated with the different values of $k$, can be successfully predicted with the proposed methodology.
@@ IV Conclusions
The main objective of this investigation was to enable effective prediction of non-linear dynamical behaviors in a dynamical system of interest. By evolving the system from given initial conditions and control parameters, the dynamics can be suitably described via recurrence plots. These plots were used to train a neural network capable of recognizing recurrence patterns and distinguishing between behaviors. In this study, two well-known non-linear maps were used to exemplify the use of this methodology. Nevertheless, the presented approach is easily generalized within a broad scope, being readily applicable to real-world phenomena. The presented results, along with the previous works on the classification of chaotic behaviors in the literature, highlight the effectiveness of predicting non-linear dynamics on various dynamical systems.
The chosen CNN architecture proved to be robust to changes in recurrence patterns from both systems, achieving relatively high accuracy in forecasting parameters for both the one-dimensional Logistic map and the two-dimensional Standard map. Notably, performance improvements were observed when focusing training on specific parameter ranges, such as the chaotic regime in the Logistic map or the pre-critical regime in the Standard map. These results suggest that further performance improvements could be achieved through larger sampling sizes and possibly the application of transfer learning strategies.
To further extend the scope of this methodology, future work could explore the use of state-of-the-art architectures such as ResNet, EfficientNet, and InceptionV3. Training these networks on a large-scale database of RPs, similar to the ImageNet project [30] could enhance the ability to identify real-world dynamical systems. This would enable the neural network to characterize multiple systems and relate them to unknown dynamical behaviors, providing a powerful tool for the classification and prediction of diverse non-linear systems.
@@ Acknowledgements.
Appendix A Optimal parameters selection
@@
The Logistic map was selected for these tests due to its ability to characterize the network’s precision across various ranges, thus capturing its well-known behaviors throughout the $r$ range. The training dataset for the neural networks initially consisted of $10^{4}$ unique RP images, generated using uniformly selected values of $r$, while the hold-out test set contained 500 examples. These sample sizes were fixed across all simulations to manage memory consumption effectively, both in terms of RAM and VRAM, as image resolution increased. The results are presented in Fig. 8.
@@
An argument can also be made for the increased complexity of the dynamics when considering higher resolutions, which result from allowing for more iterations of the maps. These more complex RPs can compromise the efficiency of the pattern recognition capability of the NNs, which is especially true for chaotic behavior as observed in classes B and C in Fig. 8. Indeed, Fig. 9 illustrates the compression of such patterns in an increasing pixel resolution and, consequently, maximum iteration time $T$.
Table 3: RMSE and the defined percentage of correct predictions $c(m)$ for the regression tasks considering the Standard map with ICs $(q_{0},p_{0})=(\delta\times 10^{-\gamma},\pi)$. Here $\delta=1.0$ and $k\in[0.1,k_{c})$.
@@
The next additional parameter that needs to be carefully selected is order $\gamma$ for the distance between the pair of ICs $(q_{0},p_{0})=(\delta\times 10^{-\gamma},\pi)$ and the unstable fixed point at $(0,\pi)$ in the Standard map. Since only chaotic dynamics is analyzed in this case, it is fair to assume that the optimal resolution would be the same as previously chosen for the Logistic map, especially considering the low RMSE for the region of chaotic behavior, namely B and C classes. With that, several simulations were performed considering the range of the characteristic control parameter $k\in[0.1,k_{c})$, a fixed $\delta=1$ and different values for the order $\gamma=5,6,7,8~{}\text{and}~{}9$. The results of RMSE and the percentage of correct predictions $c(m)$ are presented in Tab. 3.
From the comparison of the results presented in Tab. 3, $\gamma=8$ was chosen as the optimal order, meaning that the pair of ICs $(q_{0},p_{0})$ is effectively selected as $(10^{-8},\pi)$. As discussed in detail in Sec. II.2, this special pair of ICs warrants chaotic dynamics even for relatively small values of the control parameter $k$, allowing a suitable comparison between different behaviors within Hamiltonian chaos. Nevertheless, the results shown in Tab. 3 are expected, but also intriguing. On the one hand, RMSE values should indeed be inversely proportional to $\gamma$, meaning that the closer the ICs are to the unstable fixed point, the NN would be able to recognize even more different chaotic behaviors. On the other hand, $\gamma=9$ produces an unexpected outcome, since it should be the lowest value of RMSE and $c(m)$. This result is still unclear and will be addressed in future works.

### v2

@@ II.3 Convolutional Neural Networks and metrics
Each network was trained independently in two different RP train sets: the first set containing data from the logistic map generated with parameters $r\in[3.0,4.0]$, and the second with $r$ immediately after the onset of chaos; while for the standard map, $k\in[0.1,0.971635)$ and a larger range $k\in[0.1,2.0]$ were used to generate the data sets for this dynamic. The amount of samples, originated from the same sets of ICs, remained constant for the analysis focused on the onset of chaos and $k<k_{c}\approx 0.971635$, resulting in larger training sets for these regions where the chaotic behaviour is predominant.
@@
The evaluation of the performance of the trained models used the following metrics: the weighted $f_{1}$ score was employed to assess the accuracy of inferences across the defined intervals of $r$ to which the testing RP samples belonged in the logistic map. For the regression tasks, a point-wise error, based on the absolute differences between the actual and inferred parameter values, was used to evaluate the estimated $r$ and $k$ values for the logistic and standard maps, respectively.
Additionally, considering an acceptable margin $m$, a given inference can be categorized as follows
@@
where $y_{i}$ and $\hat{y_{i}}$ represent the actual and inferred parameter values for sample $i$, respectively. The Heaviside step function ensures that if $|y_{i}-\hat{y_{i}}|<m$, $C_{i}=1$, indicating that the inferred value is within the acceptable margin $m$ from the actual value; otherwise, $C_{i}=0$, representing an incorrect inference.
The total number of inferences within the $m$-tolerance is then given by $C(m)=\sum_{i=1}^{N}C_{i}(m)$, where $N$ is the total number of test samples. Furthermore, the percentage of correct inferences is defined as
@@
where, in practical terms, if $m=0.01$, the percentage $c(0.01)$ expresses how accurately the NN can, indeed, infer the parameter values up to the third decimal place.
@@ III Results and discussions
The first set of results comes from applying the aforementioned methodology to the logistic map. $4\times 10^{4}$ RPs, representing distinct non-linear dynamical behaviours from Eq. (7), were constructed considering fixed $RR=20\%$ and $T=100$. The task of the trained NNs is to infer both the value and range of the characteristic control parameter $r$, respectively through a regression and classification task. Table 1 presents the outcome of the models for the hold-out test samples.
Table 1: Results of the classification task on the logistic map, employing $f_{1}$ score as the metric; and the regression task, using RMSE and $c(m)$ defined in Eq. (11), both for the test set, and for each interval of $r$ as described in Sec. II.2.1.
@@
(A3) 0.0475 95.7% 22.9% 0.971 17
(B) 0.0206 100% 37.3% 0.988 608
(C) 0.0235 100% 0% 0.981 375
weight. avg. 0.0334 - - 0.985 1000
Figure 6: Inferred values ($\hat{r}$) compared to actual parameter ($r$) values of the test set from the logistic map in the regression task. The colour axis illustrates the dispersion on the absolute difference $|r-\hat{r}|$ in both cases: (a) $r\in[3.0,4.0]$ and; (b) $r\in[3.56,4.0]$.
Notice that the weighted accuracy obtained, shown in column “$f_{1}$ score” for training the NN almost exclusively on the chaotic region is higher than currently reported in literature [12, 13], and also improved while comparing with RPs originated from $r\in[3.0,4.0]$. As for the regression results displayed on columns “RMSE, $c(0.1)$ and $c(0.01)$”, the same improvement on characterizing chaotic dynamics is observed when focusing solely on the chaotic region, where at least $86.93\%$ of inferred parameters $\hat{r}$ are precise within two decimal places to the actual parameter $r$.
Figure 6 illustrates the regression results for the logistic map. In Fig. 6 (a), despite the presence of a distinct light-grey point far from the main regression trend, where the proposed NN inferred a value likely associated with the periodic windows within region (C), it is clear that the NN can successfully infer the parameter values, particularly before the region (B), onset of chaos at $r\approx 3.57$. Beyond this point, the inferences become less accurate as the chaotic regime dominates the dynamics. To further investigate, the same number of RP samples was used to train the NN, but considering only the range $r\in[3.56,4.0]$. The regression results for this specific range are shown in Fig. 6 (b).
In comparison to (a), Fig. 6 (b) shows an inferior agreement between the inferred and actual parameter values. This is due to the fact that the chaotic regime may present many distinct recurrence patterns, as earlier evidenced by the RPs within this range in Fig. 2. These many faces of chaotic behaviour are expected for a finite-time analysis, making the effective parameter inference more challenging, and consequently lowering the NN accuracy. Additionally, regions between $3.93.8<r<$ 3.9, where the agreement is even less accurate, correlate to the rise of periodic and quasi-periodic behaviours due to islands of stability of said region. Yet, as observed in Tab. 1, this region is successfully classified, presenting relatively high $f_{1}$ score.
It is worth mentioning that all results presented for the logistic map are further compared with those obtained from raw time-series data analysis using an equivalent NN architecture in Appendix A. These comparisons highlight that the proposed RP-based methodology provided better parameter inferences.
Figure 7: Inferred values ($\hat{k}$) compared to actual parameter ($k$) values of the test set from the standard map in the regression task. The colour axis illustrates the dispersion on the absolute difference $|k-\hat{k}|$ in both cases: (a) $k\in[0.1,k_{c})$ and; (b) $k\in[0.1,2.0)$.
Table 2: RMSE and overall precision $c(m)$ for the regression task for RPs generated from the standard map.
@@
$[0.1,k_{c})$ 0.0348 98.8% 24.0%
$[0.1,2)$ 0.0677 85.5% 11.8%
Considering the standard map, Tab. 2 summarizes the performance of the trained NN with the RPs from the different non-linear dynamical behaviours of Eqs. (8) and (9). Again, $4\times 10^{4}$ RPs were constructed considering $T=100$, but now $RR=10\%$. In contrast to the logistic map, all RPs of the standard map represent chaotic behaviours, specifically different finite-time trajectories in Hamiltonian chaos. In that sense, employing the process outlined in Sec. II.1 for a relatively high $RR$ may require longer computational simulations, so $RR=10\%$ was chosen. Additionally, since only chaotic behaviour is analysed, classification tasks were neglected.
It is important to emphasize that by selecting trajectories with specific ICs, namely $(q_{0},p_{0})=(\delta\times 10^{-8},\pi)$, where $\delta=2,4,6~{}\text{and}~{}8$, is possible to ensure chaotic behaviour for $k>0.1$. As detailed in Sec. II.2, for $k\gtrsim 0.1$, a chaotic layer surrounds the vicinity of the unstable fixed point at $(0,\pi)$ and, as $k$ increases, the phase space is filled by the expanding chaotic sea. These chaotic trajectories are suitably represented by their respective RPs as depicted earlier by Fig. 4.
Regression results for the standard map are shown in Fig. 7. Initially in (a), considering the interval $k\in[0.1,k_{c})$, the agreement between the inferred and actual parameter values is surprisingly notable for chaotic dynamics. A slight decrease in performance is observed in (b) for parameters beyond the critical value $k_{c}$. Although regions above $k>2.0$ were neglected, these results illustrate that particular finite-time chaotic behaviours, associated with the different values of $k$, can be successfully characterized via the proposed methodology.
@@ IV Conclusions
The primary aim of our work was to demonstrate a recurrence-based methodology for inferring control parameters and classifying dynamical regimes in non-linear dynamical systems. By representing the system’s trajectories through recurrence plots, we provided structured visual data that could be efficiently processed by convolutional neural networks. The CNN was trained to recognise recurrence patterns associated with distinct values of the system’s control parameters and successfully distinguished between different dynamical behaviours. Although the approach does not directly predict future system states, the deterministic nature of the analysed systems implies that the accurate inference of parameters, combined with predetermined initial conditions, enables a fitting reconstruction of their dynamical evolution. This highlights the use of RPs as suitable feature spaces for parameter inference and regime classification in non-linear dynamical systems.
The chosen CNN architecture demonstrated robustness in capturing and interpreting recurrence patterns from both studied systems, achieving good accuracy in inferring control parameters for the one-dimensional logistic map, as well as for the two-dimensional standard map. Performance was enhanced when training was focused specifically on critical dynamical regimes, such as chaotic intervals in the logistic map or the stronger chaotic regime in the standard map. These results indicate that future improvements in inference accuracy could be attained through expanded training datasets, refined sampling strategies, and potentially integrating transfer learning techniques.
Further extensions of the proposed methodology could involve exploring advanced neural network architectures, such as ResNet, EfficientNet, and InceptionV3. Developing large-scale databases of RPs, analogous to the ImageNet project [44], would likely enhance the capability to identify and characterise distinct dynamical behaviours, enabling a fitting parameter inference across a broad spectrum of real-world dynamical systems. Moreover, incorporating attention mechanisms [45] or similar interpretability-focused methods could improve understanding by revealing how neural networks prioritise features within the RPs, providing deeper insights into the fundamental characteristics of non-linear behaviours.
@@ Acknowledgements.
Appendix A Raw time-series comparison
Table 3: Results of the classification task on the time-series provided by the logistic map, employing $f_{1}$ score as the metric; and the regression task, using RMSE and $c(m)$, both for the test set for each interval of $r$ as described in Sec. II.2.1. Refer to Tab. 1 for comparison to the RP methodology.
Class RMSE $c(0.1)$ $c(0.01)$ $f_{1}$ score support
$4.03.0<r<$ 4.0
(A1) 0.1522 33.0% 0% 1.000 455
(A2) 0.0511 100% 0% 0.870 94
(A3) 0.0506 100% 0% 0.382 26
(B) 0.0600 94.3% 5.7% 0.987 246
(C) 0.0820 69.3% 6.1% 1.000 179
weight. avg. 0.1138 - - 0.962 1000
$4.03.56<r<$ 4.0
(A3) 0.0627 100% 0% 0.882 30
(B) 0.0302 100% 21.6% 0.990 575
(C) 0.0489 100% 0.5% 0.996 395
weight. avg. 0.0398 - - 0.989 1000
Figure 8: Inferred values ($\hat{r}$) compared to actual parameter ($r$) of the test set from the logistic map in the regression task, considering raw time-series under Conv1D(3,4) analysis. The color axis illustrates the absolute difference $|r-\hat{r}|$. Refer to Fig. 6 for comparison to the RP methodology.
This appendix discusses a comparison between raw time-series analysis and the RP-based methodology outlined in Sec. II, as well as the corresponding results presented in Sec. III.
As mentioned at the beginning of Sec. II.1, the construction of RPs for one-dimensional dynamical systems, such as the logistic map, is inherently tied to their time-series representation. To ensure a suitable and meaningful comparison, we conduct an analysis using the same type of CNN but applied directly to raw time-series data. The only modifications to the network architecture presented in Fig. 5 were made to accommodate the one-dimensional input, replacing the two-dimensional layers with Conv1D(3,4) and AveragePooling1D(2) for compatibility. The results of this time-series analysis are presented in Table 3.
A comparison with Table 1 shows that the time-series approach is outperformed in almost all $r$ classes, in both RMSE and the percentage of correct inferences $c$, particularly for control parameter values where chaos is observed. In addition, Fig. 8 presents the regression results for the time-series analysis. A comparison with Fig. 6 clearly illustrates that the RP-based methodology provides higher accuracy in inferring the values of the control parameters.
For the standard map, a direct comparison with the RP-based methodology is not as straightforward. Since it is a two-dimensional system, its natural phase space representation already incorporates the necessary information for recurrence analysis. While one could construct a time-series representation based on a derived quantity, such as the Euclidean distance between consecutive states, this would ultimately yield results closely related to those obtained via RP analysis. Therefore, while a strict analogy to the logistic map’s comparison is not possible, the key advantage of RP-based analysis remains evident.
Nevertheless, the results shown in Table 3 and Fig. 8 align with previous studies [32], which indicate that RPs have a greater ability to encode the overall dynamical behaviour of a system compared to raw time-series analysis, further supporting our findings.
Appendix B Optimal parameters selection
@@
The logistic map was selected for these tests due to its ability to characterize the network’s precision across various ranges, thus capturing its well-known behaviours throughout the $r$ range. The training dataset for the neural networks initially consisted of $10^{4}$ unique RP images, generated using uniformly selected values of $r$, while the hold-out test set contained 500 examples. These sample sizes were fixed across all simulations to manage memory consumption effectively, both in terms of RAM and VRAM, as image resolution increased. The results are presented in Fig. 9.
@@
An argument can also be made for the increased complexity of the dynamics when considering higher resolutions, which result from allowing for more iterations of the maps. These more complex RPs can compromise the efficiency of the pattern recognition capability of the NNs, which is especially true for chaotic behaviour as observed in classes B and C in Fig. 9. Indeed, Fig. 10 illustrates the compression of such patterns in an increasing pixel resolution and, consequently, maximum iteration time $T$.
Table 4: RMSE and the defined percentage of correct inferences $c(m)$ for the regression tasks considering the standard map with ICs $(q_{0},p_{0})=(\delta\times 10^{-\gamma},\pi)$. Here $\delta=1.0$ and $k\in[0.1,k_{c})$.
@@
The next additional parameter that needs to be carefully selected is order $\gamma$ for the distance between the pair of ICs $(q_{0},p_{0})=(\delta\times 10^{-\gamma},\pi)$ and the unstable fixed point at $(0,\pi)$ in the standard map. Since only chaotic dynamics is analyzed in this case, it is fair to assume that the optimal resolution would be the same as previously chosen for the logistic map, especially considering the low RMSE for the region of chaotic behavior, namely B and C classes. With that, several simulations were performed considering the range of the characteristic control parameter $k\in[0.1,k_{c})$, a fixed $\delta=1$ and different values for the order $\gamma=5,6,7,8~{}\text{and}~{}9$. The results of RMSE and the percentage of correct inferences $c(m)$ are presented in Tab. 4.
From the comparison of the results presented in Tab. 4, $\gamma=8$ was chosen as the optimal order, meaning that the pair of ICs $(q_{0},p_{0})$ is effectively selected as $(10^{-8},\pi)$. As discussed in detail in Sec. II.2, this special pair of ICs warrants chaotic dynamics even for relatively small values of the control parameter $k$, allowing a suitable comparison between different behaviours within Hamiltonian chaos. Nevertheless, the results shown in Tab. 4 are expected, but also intriguing. On the one hand, RMSE values should indeed be inversely proportional to $\gamma$, meaning that the closer the ICs are to the unstable fixed point, the NN would be able to recognize even more different chaotic behaviours. On the other hand, $\gamma=9$ produces an unexpected outcome, since it should be the lowest value of RMSE and $c(m)$. This result is still unclear and will be addressed in future works.